# Towards Robust Zero-Shot Reinforcement Learning

### v1

@@ 1 Introduction
Reinforcement learning (RL) has become a cornerstone of artificial intelligence, enabling transformative advances in robotics [53], autonomous systems [29], industrial control [71], and large language models (LLM) [60]. However, its real-world adoption faces two persistent challenges: the reliance on human-provided reward functions and its task-specific learning paradigm, which limits adaptability to novel or multiple tasks. These challenges sparked growing interests in zero-shot RL [57, 58, 44, 26, 4, 23, 56, 13], which enables learning a versatile agent through pretraining on reward-free transitions, and then zero-shot adapting to arbitrary reward functions during inference. This opens up new possibilities for developing general-purpose RL agents capable of generalizing across diverse tasks in open-world scenarios.
@@
Existing zero-shot RL approaches mainly fall into two categories: task/skill-conditioned RL [13, 23] and dynamic representation-based methods [4, 3, 58, 26, 56, 44]. The first category encodes demonstrations or reward functions into embeddings as conditioning signals for policy learning. This preserves the task generalization capability of policies, but often results in heavy, manually designed pretraining without optimality guarantees. The second category of methods adopt a more principled approach by decomposing the problem into dynamic representations that can be recomposed for novel tasks without retraining. Among dynamic representation-based methods, Forward-Backward representations (FB) [57, 58] have recently attracted notable attention, factorizing occupancy measures into two components: a forward representation that captures policy dynamics and a backward representation that encodes global state information. Through offline, unsupervised pretraining, the FB framework learns linearized representations that approximate value functions for arbitrary tasks, holding great promise for zero-shot generalization.
@@
However, despite the elegant theoretical framework provided by the FB representations, our empirical studies reveal that the successor measures learned through existing FB-based methods are often inconsistent and biased, thereby compromising the stability and overall performance (see Section 3 for details). In this paper, we identify two causes for the shortcomings of existing FB-based methods. First, learning complex approximators and multimodal behaviors requires highly expressive models, which current FB methods lack in both their representations and policy. Second, the offline, unsupervised pretraining stage suffers from extrapolation error due to out-of-distribution (OOD) actions, which is similar to offline RL [17, 32, 33], but exhibits more complex behavior. As our results show, naively integrating value constraints can be ineffective [26], indicating a need for a more delicate OOD regularization mechanism.
@@
Based on the above observations, we propose Behavior-REgularizEd Zero-shot RL with Expressivity enhancement (BREEZE), a novel FB-based framework that simultaneously enhances offline learning stability and zero-shot generalization capability. First, we introduce a behavior-regularized reformulation of FB, which mitigates extrapolation errors while preserving the fidelity of representations. Second, we extract the policy using a task-conditioned diffusion model, enabling high-quality multimodal action distributions in zero-shot RL settings. Finally, we employ expressive representation networks based on the attention architecture to capture complicated dynamics. We conducted extensive experiments on the ExORL [70] benchmark and the D4RL [14] Kitchen dataset, under both full datasets and small-sample data regimes. The results demonstrate that BREEZE achieves the best or near-the-best performance while exhibiting superior robustness.
@@ 2 Preliminary
Approximate dynamic programming-based RL [47] uses an action-value $Q$-function, or optionally a state-value $V$-function. The $Q$-function and $V$-function for $\pi$ starting at $s_{0},a_{0}$ under a given reward function $r:\mathcal{S}\to\mathbb{R}$ are respectively defined as $Q^{\pi}_{r}(s_{0},a_{0}):=\sum_{t\geq 0}\gamma^{t}\mathbb{E}[r(s_{t+1})|s_{0},a_{0},\pi]$ and $V^{\pi}_{r}(s_{0}):=\sum_{t\geq 0}\gamma^{t}\mathbb{E}[r(s_{t+1})|s_{0},\pi]$. The goal of the zero-shot RL problem is to train a task-agnostic policy given offline dataset $\mathcal{D}=\{s_{i},a_{i},s_{i+1}\}_{i=1}^{N}$ generated by an unknown behavior policy $\mu(\cdot|s)$, that can later generalize to any downstream task $z$ defined by a reward function $r_{\text{eval}}:\mathcal{S}\to\mathbb{R}$, i.e. find $\max_{\pi}\ \mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r_{\text{eval}}\left(s_{t+1}\right)|s_{0},a_{0},\pi]$, relying only on a small set of reward-labeled samples with no further finetuning.
@@
Given representation space $\mathbb{R}^{d}$, state distribution $\rho$, task vector $z\in\mathbb{R}^{d}$ and policy $\pi_{z}$ parameterized by $z$, FB decouples $M^{\pi_{z}}$ by a pair of forward representation $F:\mathcal{S}\times\mathcal{A}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and backward representation $B:\mathcal{S}\to\mathbb{R}^{d}$, resulting optimal policy $\pi_{z}$ as:
@@
Recent extensions of FB incorporate additional regularization to improve learning performance. For example, Jeen et al. [26] noted that the use of $\pi_{z}(s_{t+1})$ in Eq. (4) during offline learning could introduce OOD extrapolation errors. To resolve this problem, they introduce additional CQL-style [33] regularizers on $M^{\pi_{z}}$ (MCFB) or $Q_{z}$ (VCFB) in $\mathcal{L}_{\text{FB}}$ for OOD regularization (i.e., push down the corresponding values on OOD actions and push up values evaluated on offline data samples).
@@ 3 Pitfalls of Exiting FB-Based Methods
FB offers an elegant theoretical foundation for zero-shot RL; however, our empirical findings reveal that FB-based methods often produce inconsistent and biased representations in practice. To illustrate this, we evaluate the $M^{\pi_{z}}$- and $Q_{z}$-value distributions derived from the learned $F$ and $B$ representations of the vanilla FB [58], the more recent MCFB [26], and our method on ExORL [70] benchmark with RND [5] datasets. For each method, we use its $B$ representation to derive the task vector $z$ under default settings with a fixed number of transitions. The $M^{\pi_{z}}$ values are computed from episode-start state-action pairs $(s_{0},a_{0})\sim\rho_{0}$ and randomly sampled transitions $s_{+}\sim\mathcal{D}$, while $Q$-values are evaluated using policy-induced actions $a\sim\pi_{z}(s)$. Results across additional environments and with method VCFB [26] are provided in Appendix D.3.
@@
Figure 2: Investigations of the modeling architecture on the vanilla FB [58]. We utilize the peak observed in the training stage of each run as a proxy for the architecture’s capability ceiling.
@@ 4.1 Behavior-regularized Optimization
where $z$ is the task vector and $\mu(a|s)$ is the behavior policy in dataset $\mathcal{D}$. Compared with the traditional formulation of state-value function, this optimization problem aims to learn an optimal task-conditioned $V$-function solely from dataset samples without explicitly estimating $\mu(a|s)$. It can be solved through a class of value objectives according to different regularization formulations [67, 18, 30]. In this work, we leverage the commonly used expectile regression as in IQL [30]:
@@
where $L_{2}^{\tau}(u)=|\tau-\mathbb{I}(u>0)|u^{2}$, with $\tau>0.5$ serving as the expectile parameter. By minimizing the above equations, we can obtain a well-regularized and nearly optimal state-value function $V_{\pi_{z}}$ for different tasks. Given the reward function $r_{z}$, associated with task $z$, we have the following modified policy evaluation operator $\mathcal{T}^{\pi}$ given by
@@
$\mathcal{L}_{F\text{-reg}}=\mathbb{E}_{\begin{subarray}{c}(s,a,s^{\prime})\sim\mathcal{D}\end{subarray}}\left[\left(F(s,a,z)^{\top}z-B(s^{\prime})^{\top}\mathbb{E}_{D}[BB^{\top}]^{-1}z-\gamma V_{\pi_{z}}(s_{t+1},z)\right)^{2}\right].$
@@
Having introduced regularization for the value function, we now turn to policy learning, which prior work suggests can be equally or even more critical for final performance [43]. A good policy should effectively leverage the high-quality behavior and further generalize near data distribution. To ensure that the policy is well regularized by the dataset samples, we replace the policy objective in Eq. (3) with the following behavior-regularized optimization problem:
@@ 4.2 Policy Extraction via Task-Conditioned Diffusion Model
To compute the solution of the weighted behavior cloning objective in Eq. (11) in practice, a key challenge arises: how to efficiently model and sample from potentially highly complex and diverse distributions distilled from a reweighted behavior policy?
@@
4.3 Expressivity Enhancement for Representations Modeling
The effectiveness of any policy—particularly a highly expressive diffusion policy—in leveraging learned action weights hinges on accurate value estimation. Biased value estimates can degrade policy learning to simple imitation, failing to redistribute probability mass toward superior actions. To fully exploit the representational capacity of our diffusion policy, the underlying value representations must precisely capture complex task and dynamics relationships. We thus introduce enhanced network architectures for the forward ($F$) and backward ($B$) representations.
Figure 3: The architectures in BREEZE
@@ 4.3 Expressivity Enhancement for Representations Modeling
The forward network encodes state-task and state-action pairs using two separate linear encoders, following Touati et al. [58]. The state-task encoder captures task-conditioned dynamics, while the state-action encoder extracts agent behavior patterns. These two feature sets correspond to semantically distinct concepts in the MDP. To accurately approximate the future occurrence measure $M(s,a,s^{\prime},z)$ in Eq. 3, the representation $F$ must integrate information from both behavioral and task contexts, capturing their interdependencies effectively. We therefore model the two encoded features as a length-2 sequence and process them through self-attention blocks. This allows bidirectional feature refinement between task conditions and agent behaviors. The resulting representation is projected to a $d$-dimensional space via a linear layer.
@@
As defined in Eq. 3, the backward representation $B$ encodes state-level structural information, serving as a global embedding of the environment. Intuitively, $B$ should employ a more complex architecture, maintaining orthogonality and enforcing $F$ alignment. We equip $B$ with a stack of standard transformer-based networks with multi-head attention. The final output is projected to a $d$-dimensional space, consistent with the forward network.
@@
Table 1: IQM results on ExORL[70] benchamrk. We report the best overall performance when all tasks perform well. Each value is averaged over 10 rollouts on 5 random seeds.
@@
Table 3: Normalized scores on Kitchen environment
@@ 5 Experiments
Our main experiments are conducted on the ExORL [70] benchmark, which provides a variety of datasets collected by several unsupervised RL algorithms [34]. We select datasets collected by 4 algorithms: RND [5], APS [38], DIAYN [11], PROTO [69]. The experiments span 3 domains and 12 tasks (Walker: Stand, Walk, Run, Flip; Jaco: Reach Top/Bottom Left/Right; Quadruped: Stand, Walk, Run, Jump), bringing the total to 48 state-based complex tasks for performance calculation. In addition, we consider four challenging multi-stage tasks in the D4RL [14] Franka Kitchen [19] domain with two datasets (mixed and partial), which require long-horizon sequential manipulation on 4 subtasks. The overall setups on two benchmarks involve 2 locomotion domains and 2 manipulation (goal-reaching) domains. For all goal-conditioned domains, we use the backward representation to calculate the goal features as the task vector $z$.
@@
We only consider offline algorithms that could perform a zero-shot policy generalization for our evaluation. Our baselines includes: 1) SF-LAP [3, 4]: successor feature-based method with basic features to be Laplacian Eigenfunction (LAP) [64], 2) vanilla FB [57], 3) VCFB and MCFB [26]: FB-based offline algorithms with CQL [33]-style regularizer on successor measure $M^{\pi}$ and $Q$-value separately for OOD issue avoidance, 4) HILP [44]: a state-based representation modeling algorithm keeping distance-preserving temporal structures in latent space having zero-shot capability.
@@
Our experiments are designed to evaluate two key aspects: zero-shot performance and robustness. To assess the overall zero-shot capability, we experiment on the full transitions of each dataset on ExORL and D4RL Kitchen. For robustness, we first present performance curves across training time to demonstrate learning stability and then conduct experiments on a uniformly subsampled 100,000 transitions to examine performance in a limited and low-quality data coverage scenario, following the evaluation process of Jeen et al. [26]. We record the Interquartile Mean (IQM), a robust metric that reduces the influence of outliers, in the training stage. We evaluate the checkpoints every 10,000 updates with 10 rollouts for each experiment. On Kitchen, we multiply the returns by 25 for normalization, following Park et al. [44]. Further implementation and experimental details are provided in Appendix B.

### v2

@@ 1 Introduction
Reinforcement learning (RL) has become a cornerstone of artificial intelligence, enabling transformative advances in robotics [53], autonomous systems [29], industrial control [71], and large language models (LLM) [60]. However, its real-world adoption faces two persistent challenges: the reliance on human-provided reward functions and its task-specific learning paradigm, which limits adaptability to novel or multiple tasks. These challenges sparked growing interest in zero-shot RL [57, 58, 44, 26, 4, 23, 56, 13], which enables learning a versatile agent through pretraining on reward-free transitions and then zero-shot adaptation to arbitrary reward functions during inference. This opens up new possibilities for developing general-purpose RL agents capable of generalizing across diverse tasks in open-world scenarios.
@@
Existing zero-shot RL approaches mainly fall into two categories: task/skill-conditioned RL [13, 23] and dynamic representation-based methods [4, 3, 58, 26, 56, 44]. The first category encodes demonstrations or reward functions into embeddings as conditioning signals for policy learning. This preserves the task generalization capability of policies, but often results in heavy, manually designed pretraining without optimality guarantees. The second category of methods adopts a more principled approach by decomposing the problem into dynamic representations that can be recomposed for novel tasks without retraining. Among dynamic representation-based methods, Forward-Backward representations (FB) [57, 58] have recently attracted notable attention, factorizing occupancy measures into two components: a forward representation that captures policy dynamics and a backward representation that encodes global state information. Through offline, unsupervised pretraining, the FB framework learns linearized representations that approximate value functions for arbitrary tasks, holding great promise for zero-shot generalization.
@@
However, despite the elegant theoretical framework provided by the FB representations, our empirical studies reveal that the successor measures learned through existing FB-based methods are often inconsistent and biased, thereby compromising the stability and overall performance (see Section 3 for details). In this paper, we identify two causes for the shortcomings of existing FB-based methods. First, learning complex approximators and multimodal behaviors requires highly expressive models, which current FB methods lack in both their representations and policy. Second, the offline, unsupervised pretraining stage suffers from extrapolation error due to out-of-distribution (OOD) actions, which is a similar problem to offline RL [17, 32, 33], but exhibits more complex behavior. As our results show, naively integrating value constraints can be ineffective [26], indicating a need for a more delicate OOD regularization mechanism.
@@
Based on the above observations, we propose Behavior-REgularizEd Zero-shot RL with Expressivity enhancement (BREEZE), a novel FB-based framework that simultaneously enhances offline learning stability and zero-shot generalization capability. First, we introduce a behavior-regularized reformulation of FB, which mitigates extrapolation errors while preserving the fidelity of representations. Second, we extract the policy using a task-conditioned diffusion model, enabling high-quality multimodal action distributions in zero-shot RL settings. Finally, we employ expressive representation networks based on the attention architecture to capture complicated dynamics. We conducted extensive experiments on the ExORL benchmark [70] and the D4RL Kitchen dataset [14], under both full datasets and small-sample data regimes. The results demonstrate that BREEZE achieves the best or near-the-best performance while exhibiting superior robustness.
@@ 2 Preliminary
Approximate dynamic programming-based RL [47] uses an action-value $Q$-function, or optionally a state-value $V$-function. The $Q$-function and $V$-function for $\pi$ starting at $s_{0},a_{0}$ under a given reward function $r:\mathcal{S}\to\mathbb{R}$ are respectively defined as $Q^{\pi}_{r}(s_{0},a_{0}):=\sum_{t\geq 0}\gamma^{t}\mathbb{E}[r(s_{t+1})|s_{0},a_{0},\pi]$ and $V^{\pi}_{r}(s_{0}):=\sum_{t\geq 0}\gamma^{t}\mathbb{E}[r(s_{t+1})|s_{0},\pi]$. The goal of the zero-shot RL problem is to train a task-agnostic policy given an offline dataset $\mathcal{D}=\{s_{i},a_{i},s_{i+1}\}_{i=1}^{N}$ generated by an unknown behavior policy $\mu(\cdot|s)$, that can later generalize to any downstream task $z$ defined by a reward function $r_{\text{eval}}:\mathcal{S}\to\mathbb{R}$, i.e. find $\max_{\pi}\ \mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r_{\text{eval}}\left(s_{t+1}\right)|s_{0},a_{0},\pi]$, relying only on a small set of reward-labeled samples with no further finetuning.
@@
Given a representation space $\mathbb{R}^{d}$, a state distribution $\rho$, a task vector $z\in\mathbb{R}^{d}$ and a policy $\pi_{z}$ parameterized by $z$, FB expresses $M^{\pi_{z}}$ as the product of a forward representation $F:\mathcal{S}\times\mathcal{A}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and a backward representation $B:\mathcal{S}\to\mathbb{R}^{d}$, resulting in an optimal policy $\pi_{z}$:
@@
Recent extensions of FB incorporate additional regularization to improve learning performance. For example, Jeen et al. [26] noted that the use of $\pi_{z}(s_{t+1})$ in Eq. (4) during offline learning could introduce OOD extrapolation errors. To resolve this problem, they introduce additional CQL-style [33] regularizers on $M^{\pi_{z}}$ (MCFB) or $Q_{z}$ (VCFB) in $\mathcal{L}_{\text{FB}}$ for OOD regularization (i.e., decrease values for OOD actions and increase those for data samples).
@@ 3 Pitfalls of Exiting FB-Based Methods
FB offers an elegant theoretical foundation for zero-shot RL; however, our empirical findings reveal that FB-based methods often produce inconsistent and biased representations in practice. To illustrate this, we evaluate the $M^{\pi_{z}}$- and $Q_{z}$-value distributions derived from the learned $F$ and $B$ representations of the vanilla FB [58], the more recent MCFB [26], and our method on ExORL benchmark [70] with RND [5] datasets. For each method, we use its $B$ representation to derive the task vector $z$ under default settings with a fixed number of transitions. The $M^{\pi_{z}}$ values are computed from episode-start state-action pairs $(s_{0},a_{0})\sim\rho_{0}$ and randomly sampled batch of transitions $s_{+}\sim\mathcal{D}$, while $Q$-values are evaluated using policy-induced actions $a\sim\pi_{z}(s)$. Results across additional environments and with method VCFB [26] are provided in Appendix D.3.
@@
Figure 2: Investigation of the modeling architecture on the vanilla FB [58]. We utilize the peak observed in the training stage of each run as a proxy for the architecture’s capability ceiling.
@@ 4.1 Behavior-regularized Optimization
where $z$ is the task vector and $\mu(a|s)$ is the behavior policy in the dataset $\mathcal{D}$. Compared with the traditional formulation of the state-value function, this optimization problem aims to learn an optimal task-conditioned $V$-function solely from dataset samples without explicitly estimating $\mu(a|s)$. It can be solved through a class of value objectives according to different regularization formulations [67, 18, 30]. In this work, we leverage the commonly used expectile regression as in IQL [30]:
@@
where $L_{2}^{\tau}(u)=|\tau-\mathbb{I}(u<0)|u^{2}$, with $\tau>0.5$ serving as the expectile parameter. By minimizing the above equations, we can obtain a well-regularized and nearly optimal state-value function $V_{\pi_{z}}$ for different tasks. Given the reward function $r_{z}$, associated with task $z$, we have the following modified policy evaluation operator $\mathcal{T}^{\pi}$ given by
@@
$\mathcal{L}_{F\text{-reg}}=\mathbb{E}_{\begin{subarray}{c}(s,a,s^{\prime})\sim\mathcal{D}\end{subarray}}\left[\left(F(s,a,z)^{\top}z-B(s^{\prime})^{\top}\mathbb{E}_{D}[BB^{\top}]^{-1}z-\gamma V_{\pi_{z}}(s^{\prime},z)\right)^{2}\right].$
@@
Having introduced regularization for the value function, we now turn to policy learning, which prior work suggests can be equally or even more critical for final performance [43]. A good policy should effectively leverage high-quality behavior and further generalize near data distribution. To ensure that the policy is well regularized by the dataset samples, we replace the policy objective in Eq. (3) with the following behavior-regularized optimization problem:
@@ 4.2 Policy Extraction via Task-Conditioned Diffusion Model
To compute the solution of the weighted behavior cloning objective in Eq. (11) in practice, a key challenge arises: how to efficiently model and sample from potentially highly complex and diverse distributions distilled from a re-weighted behavior policy?
@@
4.3 Expressivity Enhancement for Representation Modeling
The effectiveness of any policy—particularly a highly expressive diffusion policy—in leveraging learned action weights hinges on accurate value estimation. Biased value estimates can degrade policy learning to simple imitation, failing to redistribute probability mass toward superior actions. To fully exploit the representational capacity of our diffusion policy, the underlying value representations must precisely capture complex task and dynamic relationships. We thus introduce enhanced network architectures for the forward ($F$) and backward ($B$) representations.
Figure 3: Architecture of BREEZE
@@ 4.3 Expressivity Enhancement for Representation Modeling
The forward network encodes state-task and state-action pairs using two separate linear encoders, following Touati et al. Touati et al. [58]. The state-task encoder captures task-conditioned dynamics, while the state-action encoder extracts agent behavior patterns. These two feature sets correspond to semantically distinct concepts in the MDP. To accurately approximate the measure $M(s,a,s^{\prime},z)$ in Eq. 3, the representation $F$ must integrate information from both behavioral and task contexts, capturing their interdependencies effectively. We therefore model the two encoded features as a length-2 embedding sequence and process them through self-attention blocks. This allows bidirectional feature refinement between task conditions and agent behaviors. The resulting representation is projected onto a $d$-dimensional space via linear layers.
@@
As defined in Eq. 3, the backward representation $B$ encodes state-level structural information, serving as a global embedding of the environment. Intuitively, $B$ should employ a more complex architecture, maintaining orthogonality and enforcing $F$ alignment. We equip $B$ with a stack of standard transformer-based networks with multi-head attention. The final output is projected onto a $d$-dimensional space, consistent with the forward network.
@@
Table 1: IQM results on ExORL benchmark. We report the best overall performance when all tasks perform well. Each value is averaged over 10 rollouts on 5 random seeds.
@@
Figure 4: Normalized scores on Kitchen environment
@@ 5 Experiments
Our main experiments are conducted on the ExORL benchmark [70], which provides a variety of datasets collected by several unsupervised RL algorithms [34]. We select datasets collected by 4 algorithms: RND [5], APS [38], DIAYN [11], and PROTO [69]. The experiments span 3 domains and 12 tasks (Walker: Stand, Walk, Run, Flip; Jaco: Reach Top/Bottom Left/Right; Quadruped: Stand, Walk, Run, Jump), bringing the total to 48 state-based complex tasks for performance calculation. In addition, we consider four challenging multi-stage tasks in the D4RL [14] Franka Kitchen domain [19] with two datasets (mixed and partial), which require long-horizon sequential manipulation on 4 subtasks. The overall setup involves 2 locomotion domains and 2 manipulation (goal-reaching) domains. For all goal-conditioned domains, we use the backward representation to calculate the goal features as the task vector $z$.
@@
We only consider offline algorithms that could perform a zero-shot policy generalization for our evaluation. Our baselines include: 1) SF-LAP [3, 4]: successor feature-based method with basic features to be Laplacian Eigenfunction (LAP) [64], 2) vanilla FB [57], 3) VCFB and MCFB [26]: FB-based offline algorithms with CQL-style regularizer [33] on successor measure $M^{\pi}$ and $Q$-value separately for OOD issue avoidance, 4) HILP [44]: a state-based representation modeling algorithm keeping distance-preserving temporal structures in latent space having zero-shot capability.
@@
Our experiments are designed to evaluate two key aspects: zero-shot performance and robustness. To assess the overall zero-shot capability, we experiment on the full transitions of each dataset on ExORL and D4RL Kitchen. For robustness, we first present performance curves across training time to demonstrate learning stability and then conduct experiments on a uniformly subsampled 100,000 transitions to examine performance in a limited and low-quality data coverage scenario, following the evaluation process of Jeen et al. [26]. We record the Interquartile Mean (IQM), a robust metric that reduces the influence of outliers. We evaluate the checkpoints every 10,000 updates with 10 rollouts for each experiment. On Kitchen, we multiply the returns by 25 for normalization, following Park et al. [44]. Further implementation and experimental details are provided in Appendix B.